# HiMo: High-Speed Objects Motion Compensation in Point Clouds

### v2

@@ VI-D Computational Cost Comparison
Figure 12: Impact of motion distortion on 3D detection using TransFusion [56] applied to Scania highway scenes. The examples are from the same scene at different timestamps during a 2-second sequence ($t=5.2$, $6.2$, $7.2$ s). Top: detection results on data with ego-motion compensation show three different failure modes. (1.a) shows an over-extended bounding box with incorrect orientation; (2.a) shows missing geometry due to distortion; (3.a) shows a failure in detection (no bounding box). Bottom: detection results on undistorted data processed by HiMo pipeline. Note how the proposed data undistortion mitigates all three detection errors over the same 2s sequence. As shown in the figure, detection using HiMo-processed data produces compact, accurate, and consistent bounding boxes.
@@ Acknowledgement
Thanks to Bogdan Timus and Magnus Granström from Scania and Ci Li and Yixi Cai from KTH RPL, who helped with this work. This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation and Prosense (2020-02963) funded by Vinnova. The computations were enabled by the supercomputing resource Berzelius provided by National Supercomputer Centre at Linköping University and the Knut and Alice Wallenberg Foundation, Sweden.
@@
While several optimization-based methods (e.g., NSFP, EulerFlow, Floxels) achieve high accuracy, their runtimes span from several minutes to hours per sequence, making them unsuitable for real-time applications. In contrast, SeFlow++ offers a favorable trade-off between accuracy and efficiency, completing each sequence in just 8.2 seconds. Notably, SeFlow++ remains competitive in both performance and efficiency even when compared to recent entries (e.g., SemanticFlow, Floxels, and VoteFlow).

### v3

@@ VI-C Qualitative Results
Figure 12: Impact of motion distortion on 3D detection using TransFusion [56] applied to Scania highway scenes. The examples are from the same scene at different timestamps during a 2-second sequence ($t=5.2$, $6.2$, $7.2$ s). Top: detection results on data with ego-motion compensation show three different failure modes. (1.a) shows an over-extended bounding box with incorrect orientation; (2.a) shows missing geometry due to distortion; (3.a) shows a failure in detection (no bounding box). Bottom: detection results on undistorted data processed by the HiMo pipeline. Note how the proposed data undistortion mitigates all three detection errors over the same 2s sequence. As shown in the figure, detection using HiMo-processed data produces compact, accurate, and consistent bounding boxes.
@@ Acknowledgement
Thanks to Bogdan Timus and Magnus Granström from Scania and Ci Li from KTH RPL, who helped with this work. We also thank Yixi Cai, Yuxuan Liu, Peizheng Li and Shenghai Yuan for helpful discussions during revision. We thank the anonymous reviewers and the associate editor for their useful comments. The computations were enabled by the supercomputing resource Berzelius provided by National Supercomputer Centre at Linköping University and the Knut and Alice Wallenberg Foundation, Sweden.
@@
While several optimization-based methods (e.g., NSFP, EulerFlow, Floxels) achieve high accuracy, their runtimes span from several minutes to hours per sequence, making them unsuitable for real-time applications. In contrast, SeFlow++ offers a favorable trade-off between accuracy and efficiency, completing each sequence in just 8.2 seconds.